### v1

# MAPO: Advancing Multilingual Reasoning through Multilingual Alignment-as-Preference Optimization

Abstract
Though reasoning abilities are considered language-agnostic, existing LLMs exhibit inconsistent reasoning abilities across different languages, e.g., reasoning in a pivot language is superior to other languages due to the imbalance of multilingual training data. To enhance reasoning abilities in non-pivot languages, we propose an alignment-as-preference optimization framework. Specifically, we adopt an open-source translation model to estimate the consistency between answers in non-pivot and pivot languages. We further adopt the answer consistency as the preference for DPO or PPO thus optimizing the lesser reasoning. Experiments show that our method significantly improves the model’s multilingual reasoning, with better reasoning consistency across languages. Our framework achieved a $13.7\%$ accuracy improvement on out-of-domain datasets MSVAMP while preserving the competitive performance on MGSM. Moreover, we find that iterative DPO is helpful for further alignment and improvement of the model’s multilingual mathematical reasoning ability, further pushing the improvement to $16.7\%$  [^1] The project will be available at https://github.com/NJUNLP/MAPO.
MAPO: Advancing Multilingual Reasoning through Multilingual Alignment-as-Preference Optimization
Shuaijie She , Shujian Huangthanks: Corresponding author , Wei Zou , Wenhao Zhu Xiang Liu, Xiang Geng, Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University shesj@smail.nju.edu.cn, huangsj@nju.edu.cn {zouw,zhuwh,liuxiang,gx}@smail.nju.edu.cn, chenjj@nju.edu.cn
1 Introduction
The reasoning ability of large-scale models has attracted much attention at present. Existing work improves the reasoning ability of models by constructing a large number of English fine-tuned data with corresponding inference processes Yu et al. (2023). Though, we consider reasoning to be language agnostic, existing studies Chen et al. (2023) show that due to the imbalance of pre-training and fine-tuning data across languages, the mathematical reasoning ability of existing large-scale models also varies across languages, e.g., English is far superior to that of the other languages.
To address the issue of insufficient reasoning ability in other languages,  Chen et al., 2023 translates the English questions with corresponding reasoning processes into other languages for supervised fine-tuning (Ouyang et al., 2022, SFT), thus improving the multilingual reasoning of the model. However, SFT is restricted by insufficient high-quality and diverse references with detailed reasoning in other languages, where the out-of-domain performance of the fine-tuned model is also hard to guarantee Ouyang et al. (2022). Furthermore, SFT also suffers from potential translation errors and catastrophic forgetting. Studies like Huang et al. (2023); Qin et al. (2023) circumvents LLM training by designing a two-stage prompting, which first requires the model to translate input problems from other languages into English for solutions. However, this strategy requires LLMs with proficiency in both translation and reasoning, which is considered even harder.
Intuitively, the reasoning process shall remain constant across languages, thus we can leverage the superior reasoning abilities in a pivot language, such as English, by aligning the reasoning process of non-pivot languages to the pivot. That is, we encourage problem-solving in non-pivot language to reason similar to that in the pivot language. Therefore, we propose a multilingual alignment-as-preference optimization (MAPO) framework that can improve the model’s multilingual reasoning ability with better generalization without annotating multilingual reasoning processes.
Specifically, our framework consists of two parts: preference estimation via multilingual alignment, and preference optimization. During preference estimation, we collect the reasoning processes of different languages to the same question. Instead of utilizing direct translations, we adopt a well-trained translation model to yield the generation probability of the answers in pivot language given answers in non-pivot languages. Since the higher generation probability indicates a more consistent reasoning aligned with the pivot language, the corresponding answers in the non-pivot language are favored and shall be promoted.
Finally, we adopt state-of-the-art preference optimization methods, PPO and DPO, to optimize the alignment as preference. Experiments show that MAPO effectively improves reasoning consistency across languages, i.e., improves the lesser reasoning ability in non-pivot language, achieving a 13.7% improvement in accuracy on out-of-domain datasets MVSVAMP. MAPO also retains the original reasoning ability while generalizing to out-of-domain tests compared to the SFT baseline. Moreover, MAPO can further improve the multilingual reasoning of the model by iterative DPO, outperforming all baselines including ChatGPT OpenAI (2022) across all three benchmarks.
2 Related Work
2.1 Multilingual Reasoning
A straightforward measurement to evaluate the efficacy of large language models (LLMs) lies in their proficiency in tackling complex reasoning, e.g., their performance in solving mathematical reasoning problems. Recent work Wei et al. (2023); Wang et al. (2022) has verified a substantial improvement when LLMs are guided through a step-by-step reasoning process, instead of conducting a direct answer. Efforts to refine prompting strategies Wang et al. (2022); Yao et al. (2023); Lee and Kim (2023) or to fine-tune models using datasets annotated with explicit reasoning steps Yu et al. (2023); Yue et al. (2023) have yielded promising outcomes. Some research has introduced mathematical reasoning datasets in the form of application problems, such as GSM8K Cobbe et al. (2021), NumGLUE Mishra et al. (2022), and SVAMP Patel et al. (2021).
To evaluate the multilingual reasoning capabilities of LLMs, Shi et al., 2022 proposes MGSM by manually translating 250 samples of the GSM8K test set from English into nine other languages. Subsequent researches focus on enhancing the multilingual reasoning of LLMs. Chen et al., 2023 translates the GSM8K training data into the same nine languages for supervised fine-tuning (SFT), which improves the model’s multilingual reasoning capabilities. However, SFT suffers from data scarcity and catastrophic interference. Its out-of-domain generalization is also hard to guarantee. Multilingual reasoning via LLMs remains an open challenge.
2.2 Preference Optimization
Though pretrained language models on vast training data achieve outstanding performance amongst various tasks, they still suffer from potentially harmful behaviors such as fabricating facts and offensive replies. RLHF Zheng et al. (2023); Bai et al. (2022b) further rectify these model behaviors by reinforcement learning via the preference of human feedback. RLHF introduces a reward model $r_{\theta}(x,y)$ that captures the preference nuance from the human feedback. Then, $r_{\theta}(x,y)$ scores arbitrary LLM outputs $y$ given input $x$ for iterative policy rectifications during proximal policy optimization (Schulman et al., 2017, PPO). The tuning is guided by Eq 1 to maximize the expected rewards of the LLM policy $\pi_{\theta}$ with the minimum deviation from the SFT policy:
$\mathcal{L}_{\text{ppo}}=\mathbb{E}_{(x,y)\sim D_{\pi}}[r_{\theta}(x,y)-\gamma\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{SFT}}(y|x)}],$ (1)
where $\pi_{\text{SFT}}$ is the original LLM via SFT, $\gamma$ is a hyperparameter that constrains policy updates.
Though RLHF via PPO is effective in adapting LLMs to versatile human preferences, it involves four sub-models, making the training complex and costly. DPO Rafailov et al. (2023) proposes to distill a referential SFT policy $\pi_{\text{ref}}$ by polarizing the preference. DPO tuning involves a pair of outputs $(y_{w},y_{l})$ in Eq 2:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right],$ (2)
where $y_{w}$ is favored over $y_{l}$, and $\beta$ is a hyperparameter.
3 Method
Figure 1: Illustration of our alignment framework. For brevity, we only show three sampling results and simplify the optimization processes of DPO and PPO.
We propose a framework for multilingual alignment-as-preference optimization (MAPO). Specifically, our framework includes two parts: alignment as preference estimation and preference optimization.
3.1 Alignment as Preference Estimation
Intuitively, reasoning ability is language-agnostic. However, LLM reasoning varies across different languages, where we consider a dominant language provides the upper bound in reasoning for lesser languages. Therefore, a straightforward approach is to align reasoning in non-pivot languages to a dominant pivot language. That is, we can rectify LLM’s output reasoning in non-pivot languages by favoring those answers more aligned to the pivot language.
To circumvent potential errors introduced by direct translation evaluation, we simply adopt a well-trained multilingual translation model to calculate the ‘alignment’ scores between answers in non-pivot and pivot languages as a preference. Since the translation model is optimized on a large scale parallel data in source and target languages $(\mathbf{Y},\mathbf{\bar{Y}})$ by maximizing conditional generation probability in Eq 3:
$\mathop{\arg\max}\limits_{\theta}P(\mathbf{\bar{Y}}|\mathbf{{Y}};\theta),$ (3)
with more aligned target language $\mathbf{\bar{Y}}$ to $\mathbf{Y}$, the higher the conditional generation probability $P(\bar{Y}|Y)$. Then, we input the answers in non-pivot languages and force-decode a corresponding answer in pivot language for the ‘alignment’ score. Correspondingly, answers with higher ‘alignment’ scores in non-pivot languages are preferred during the following preference optimizations.
3.2 Alignment as Preference Optimization
With ‘alignment’ scores as a preference, we adopt two state-of-the-art preference optimizations: PPO and DPO, to optimize the alignment scores of reasoning in non-pivot languages given that in a dominant pivot language. To make life easier, since English is dominant in both reasoning and translation, we adopt English as the pivot language in the rest of the paper.
3.2.1 Optimization with PPO
In our setup, the ‘alignment’ score $P(\bar{Y}|Y)$ given query $x$ depicts the preference as the reward model $r_{\theta}(x,Y)$ do in PPO, thus we directly adopt the multilingual machine translation model as the reward model. During each optimization, we collect a batch of non-English outputs $\{Y_{i}\}$ with an English output $\bar{Y}$, respectively. PPO then maximizes the expected alignment score within the batch by Eq 4:
$\mathcal{L}_{\text{ppo}}=\mathbb{E}_{i}[P(\bar{Y}|Y_{i},x)-\gamma\log\frac{\pi_{\theta}(Y_{i}|x)}{\pi_{\text{SFT}}(Y_{i}|x)}]$ (4)
3.2.2 Optimization with DPO
DPO involves an input pair $(Y_{w},Y_{l})$, where $Y_{w}$ is favored over $Y_{l}$. Correspondingly, we collect $n$ outputs within one non-English language, making $\binom{n}{2}$ pairs of $(Y_{w},Y_{l})$ ranked by the alignment scores given the same English answer $\bar{Y}$. We additionally copy and freeze the original supervised fine-tuned LLM as the reference policy $\pi_{\text{ref}}$ in Eq 2. Finally, our DPO is optimized by Eq 5:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,\bar{Y},Y_{w},Y_{l})\sim\mathcal{D}}$
$\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(Y_{w}\mid x)}{\pi_{\text{ref}}(Y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(Y_{l}\mid x)}{\pi_{\text{ref}}(Y_{l}\mid x)}\right)\right],$ (5)
3.2.3 Iterative DPO
Recent research such as LLaMA2 Touvron et al. (2023) and Anthropic Bai et al. (2022a) suggests that preference data updated through multiple iterative rounds can help synchronize the reward model and the policy model, which is beneficial for the performance of preference optimization.
Therefore, we further conduct Iterative DPO to optimize reasoning alignments across languages. Specifically, assume that initially when we train $\pi_{\theta_{i}}$ on preference data sampled by itself, we obtain $\pi_{\theta_{(i+1)}}$. Subsequently, we continue to sample based on $\pi_{\theta_{(i+1)}}$ and update it further on the newly sampled preference data.
4 Experiment
Model Bn Th Sw Ja Zh Ru De Es Fr Avg En
ChatGPT Zero-shot 33.7 42.9 46.2 45.6 46.2 48.9 50.4 50.7 50.9 46.2 50.5
MathOctopusP 13B 35.2 46.8 42.8 43.2 48.8 47.6 44.4 48.0 48.4 45.0 53.2
base model: MetaMathOctopusP 7B
Base Model 29.9 40.8 44.3 44.0 47.9 48.4 51.2 52.4 50.1 45.4 53.8
DPO 42.5 52.1 53.2 59.1 54.2 55.7 56.5 58.8 58.5 54.5 61.2
base model: MathOctopusP 7B
Base Model 27.7 35.9 39.4 41.6 42.7 44.2 44.0 45.1 45.3 40.7 46.4
MultiLingual-RFT 37.9 46.4 46.4 49.6 50.8 50.4 50.7 51.6 53.4 48.6 49.4
PPO LoRA 38.9 47.5 47.1 51.0 51.7 51.1 50.3 51.6 51.4 49.0 52.7
DPO LoRA 39.4 44.7 48.9 50.8 53.7 53.2 53.3 54.2 55.7 50.4 56.5
DPO 42.0 53.2 52.7 54.7 56.4 56.9 55.7 58.5 59.3 54.4 59.6
DPO Iter1 45.3 53.9 53.8 56.8 58.1 56.6 58.7 59.1 58.5 55.6 60.0
DPO Iter2 48.8 55.2 56.0 60.3 58.8 58.3 58.1 59.7 60.8 57.3 58.4
Table 1: Model Performances on MSVAMP test set. “Avg” represents the average performance on nine non-English languages and bold text denotes the best results of the same backbone.
Figure 2: PPL across nine languages on MSVAMP. Lower PPL indicates higher consistency in reasoning processes.
Figure 2: PPL across nine languages on MSVAMP. Lower PPL indicates higher consistency in reasoning processes.
Figure 3: ARC across nine languages on MSVAMP. Higher ARC indicate higher answer consistency.
4.1 Experiment Setup22footnotemark: 2
33footnotetext: More training details can refer to Appendix A.1
Base Models: Our experiments includes two base models: MathOctopusP-7B and MetaMathOctopusP-7B.  Chen et al. (2023) propose GSM8KInstruct by translating GSM8K Cobbe et al. (2021) to nine non-English languages. Based on the dataset, Chen et al. (2023) fine-tune the LLaMA2-7B Touvron et al. (2023), resulting in . To verify the effectiveness of our method in aligning models with stronger English reasoning abilities, we choose MetaMath-7B Yu et al. (2023) and continue to train it on GSM8KInstruct as another base model.
Alignment Datasets: NumGLUE Mishra et al. (2022) is a multi-task arithmetic reasoning benchmark. To obtain the mathematical world problems, we selected tasks 1, 4, and 8 from eight tasks as our dataset (NumGLUESub). Subsequently, we used ChatGPT to translate the questions into the same nine languages as GSM8KInstruct. NumGLUESub only has 1700 samples which is economical compared to GSM8KInstruct.
When constructing the preference dataset for DPO, we use MathOctopusP-7B as the sampling model and calculate the alignment feedback using NLLB-600M-distilled  [^4] https://huggingface.co/facebook/nllb-200-distilled-600M. We froze the parameters of both models and used them solely for inference. After that, the alignment dataset will be used for optimizing two base models respectively.
Evaluation Datasets:
  • MGSM (In domain) was gathered from a select portion of the GSM8K test set, after which native speaker rendered this subset into nine additional languages. Every linguistic variant is comprised of 250 individual test instances.
  • MNumGLUESub (In domain) We use ChatGPT to translate the test set of NumGLUESub to the same nine languages as MGSM which consists of 5530 samples in total.
  • MSVAMP (Out-of domain), obtained by translating SVAMP to the same nine language by Google Translate System and verified by three professional annotators. The dataset contains 1000 questions per language.
Evaluation Metric:
  • Accuracy: We use the accuracy of problem-solving to measure the model’s reasoning ability, with a higher accuracy representing stronger reasoning ability.
  • PPL (PPL-based Alignment): We apply a translation model (NLLB-600M-distilled) to get the perplexity of the English answer given the non-English answer. Less perplexity indicates better alignment between the reasoning processes.
  • ACR (Answer Consistency Ratio) : Let $m$ denote the set of items answered correctly in English, and $n$ denote those answered correctly in non-English. ACR is then calculated as: $\text{ACR}=\frac{|m\cap n|}{|n|}$. Higher ACR indicates a greater degree of overlap in reasoning capabilities between non-English and English languages.
Baselines: A strong baseline is the Base Model, which has acquired a degree of multilingual reasoning proficiency through fine-tuning on GSM8KInstruct. Moreover, motivated by rejection sampling techniques Yuan et al. (2023), we propose another strong baseline MultiLingual-RFT. We keep the reasoning solution which yield correct answers during sampling and then use these processes as labels to train the base model. To alleviate catastrophic forgetting, we fine-tune the model for only one epoch with lower learning rate (1e-5).
Model Bn Th Sw Ja Zh Ru De Es Fr Avg En
ChatGPT Zero-shot 31.2 38.0 40.0 36.0 44.0 43.2 46.0 47.2 41.6 40.8 54.4
base model: MetaMathOctopusP
Base Model 28.8 34.0 36.4 34.0 42.8 42.0 42.0 44.8 41.2 38.4 52.4
DPO 28.8 34.8 37.6 41.2 47.2 44.0 46.8 43.6 38.0 40.2 53.6
base model: MathOctopusP
Base Model 29.2 33.6 36.4 35.2 39.2 38.8 44.8 42.4 43.2 38.1 52.0
MultiLingual-RFT 25.6 31.2 28.8 34.0 39.2 36.0 34.8 34.4 36.4 33.4 43.2
PPO LoRA 31.2 38.4 38.4 37.2 43.6 35.2 46.0 44.0 38.8 39.2 51.2
PPO LoRA 26.8 36.0 38.4 36.8 41.6 38.4 42.8 44.4 42.8 38.7 47.2
DPO 26.8 36.4 38.4 34.4 46.8 38.4 40.4 44.4 39.2 38.4 46.8
DPO Iter1 30.4 36.0 37.6 38.0 45.2 39.6 42.0 47.6 41.2 39.7 45.2
DPO Iter2 30.8 38.0 37.6 45.2 47.2 42.0 45.2 43.2 40.8 41.1 45.6
Table 2: Model Performances on MGSM test set. “Avg” represents the average performance on nine non-English languages and bold text denotes the best results of the same backbone.
Model Bn Th Sw Ja Zh Ru De Es Fr Avg En
ChatGPT Zero-shot 36.2 42.6 47.2 58.1 60.6 42.6 41.5 54.9 39.4 47.0 70.6
base model: MetaMathOctopusP
Base Model 24.5 35.2 36.2 44.1 45.8 39.2 32.6 45.2 36.7 37.7 52.5
DPO 37.3 42.0 41.6 51.6 54.8 44.4 48.2 52.4 45.2 46.4 64.0
base model: MathOctopusP
Base Model 26.6 30.9 34.3 40.9 44.4 36.0 32.6 42.0 36.2 36.0 46.9
MultiLingual-RFT 38.0 42.9 41.8 48.2 51.6 45.2 42.9 49.3 42.9 44.8 51.2
PPO LoRA 34.3 41.1 40.7 45.6 49.3 39.5 34.3 46.7 36.3 40.9 51.6
DPO LoRA 27.9 31.8 36.5 39.2 42.0 39.7 37.1 41.8 36.3 36.9 46.7
DPO 32.6 41.4 42.0 42.7 45.0 42.2 40.7 47.8 40.9 41.7 47.8
DPO Iter1 35.6 39.7 42.2 45.6 50.7 44.1 42.6 49.3 42.9 43.6 49.7
DPO Iter2 41.8 45.8 46.9 52.9 54.4 49.9 50.7 54.0 51.4 49.8 55.9
Table 3: Model Performances on MNumGLUESub test set. “Avg” represents the average performance on nine non-English languages and bold text denotes the best results of the same backbone.
5 Experiment Results
5.1 Preference Optimization Advances Multi-lingual Reasoning Ability Effectively
Based on the results in both three Table 1,2,3, our method has effectively enhanced the reasoning capabilities of the base model. Across the average non-English reasoning accuracy rates on three datasets, our approach has outperformed ChatGPT and achieved state-of-the-art performance. Specifically, even on the test set MGSM, which corresponds to the base model fine-tuning data, our method still effectively improved the reasoning accuracy by 3%. This enhancement effect is even more significant on the out-of-domain dataset MSVAMP, where we achieved an average accuracy improvement of 16.6%. Meanwhile, MultiLingual-RFT, using the same sampling results, only get an improvement of 8.7%. This demonstrates the efficiency of our method.
From the perspective of individual languages, all languages have experienced an increase in capabilities through alignment. The effect is relatively significant on some low-resource languages that previously underperformed. Our method has increased the accuracy on MSVAMP for Bengali, Thai, Swahili, and Japanese by 21.1%, 19.3%, 16.6%, and 18.7%, respectively. Moreover, the medium-to-high resource languages, which originally performed relatively well on the base model, have also continuously gained improvements in capabilities as the alignment progresses.
It’s surprising that our training data does not contain English, but the accuracy on English questions has been greatly improved. We believe this is primarily due to the alignment facilitating a more consistent understanding of reasoning across different languages, which contributes to the enhancement of its capabilities, including performance in English. Besides, our method also achieved considerable improvements on the MetaMathOctopusP 7B, which has stronger English reasoning capabilities, proving the robustness of our approach across different base models.
5.2 Alignment is the Key to Enhanced Multilingual Reasoning
We use PPL and ACR to evaluate the degree of alignment in reasoning processes and final answers between other languages and English, respectively.
Regarding the alignment of the reasoning process, as shown in Figure 3, our method effectively improves the consistency of the reasoning process, particularly for languages where the base model and ChatGPT had poorer alignment, such as Bengali, Thai, and Japanese. This proves that the model can show more similar reasoning thinking to English on non-English after alignment.
Additionally, we notice that alignment also contributes to more consistent final answers. A higher ARC indicates that there is a greater overlap between the questions answered correctly in non-English and those answered correctly in English. From the results in Figure 3, our alignment framework has greatly increased the ARC for each language. This means that the performance gains in Table 1 stem from the parts that intersect with English. These observations demonstrate that our method effectively aligns the model’s non-English reasoning processes with English, thereby enhancing reasoning abilities in non-English languages.
5.3 Preference Optimization Brings a More Generalizable Improvement in Multilingual Reasoning Ability
As indicated in Table 3, both SFT and DPO have achieved performance improvements on the test set of the MNumGLUESub. However, on the out-of-domain test set MSVAMP, DPO achieved a 13.7% increase in accuracy, which is significantly higher than the 7.9% increase achieved by SFT. More surprisingly, while DPO effectively improving the multilingual reasoning performance on both two datasets, it also maintained competitive reasoning performance on MGSM. Iterative DPO could even lead to further enhancements in performance. In contrast, SFT exhibited a 4.7% performance degradation.
We believe that this is mainly because model directly learns the given labels during SFT. The labels inevitably have distributional characteristics unique to the dataset, which are difficult to generalize to other datasets. Conversely, our method does not give the model "correct answers", but rather guides the model to generate outputs that are more aligned with the alignment characteristics in the space based on preferences, which is more effective in allowing the model to master the multi-language alignment learning objective, and this preference can better generalize to out-of-domain data sets.
6 Analysis
GSM-INS MNUM-Sub
Valid Sample 575 895
En Self-Bleu 0.81 0.53
Non-En Self-Bleu 0.84 0.60
Avg Acc on MSVAMP 31.6 54.4
Table 4: Comparison between GSM8KINSTRUCT (GSM-INS) and MNumGLUESub(MNUM-Sub)
English Question
182 students suggested adding mashed potatoes while others suggested adding bacon to the menu . if 166 more students suggested adding bacon than those that suggested mashed potatoes how many students suggested bacon ?
Chinese Question
182名学生建议在菜单中加入土豆泥,而其他人建议加入培根。如果有比建议加入土豆泥的学生多166人建议加入培根,那么有多少学生建议加入培根?
English Solution
The number of students who suggested mashed potatoes is 182. The number of students who suggested bacon is 182+166=348.
Chinese Solution [Before Alignment]
建议加入土豆泥的学生比建议加入培根的学生多166人,所以两边减去166得到差值为182 - 166 = 16,因此,建议加入培根的学生人数为16。(There are 166 more students recommending mashed potatoes than bacon so subtract 166 from both sides to get a difference of 182-166 = 16. Therefore, the number of students recommended to join Bacon is 16.)
[PPL]: 2.65
[After Alignment]
建议加入土豆泥的学生人数是182。建议加入培根的学生人数是182 + 166 = 348。 (The number of students who suggested mashed potatoes is 182. The number of students who suggested bacon is 182+166=348.)
[PPL]: 0.97
Table 5: An example from the MSVAMP test set. The example clearly demonstrates that, through our preference optimization method, the model successfully corrects erroneous reasoning processes and generates reasoning processes that is more consistent with English.
6.1 Why Not Sample on SFT Dataset
A natural question is whether we can simply sample and optimize preference on the SFT dataset (like GSM8KINSTRUCT). Experimental results indicate that the SFT dataset is not suitable as a sampling dataset for preference optimization. We randomly choose 1000 samples from the sampling results of both the GSM8KINSTRUCT and MNumGLUESub datasets. To ensure fairness, we define the samples with more than 5 answers after deduplication as valid sample and calculated the self-bleu Zhu et al. (2018) on first 5 answers within the samples (self-bleu calculates the overlap between a group of text contents, which is widely used to measure the diversity of generated results).
The analysis results are shown in the Table 4. The valid sample collected on the SFT dataset is significantly reduced, and the self-bleu is significantly higher than that of other datasets. This effectively proves that SFT will make the model quickly remember the given standard answer, which leads to a serious degradation in the diversity of the generation. The policy model struggles to learn the alignment preferences we truly hope to optimize and it easily falls into the trap of overfitting. Fortunately, compared to complex reasoning processes, multilingual questions and instructions are relatively easy to obtain, and their construction costs are lower. We only use MNumGLUE which has only 1700 samples, and the prospect of sampling based on more multilingual questions and optimizing for alignment preferences remains for future research.
Figure 4: Average accuracy on Non-English math problems and ACR versus preference optimization step on MSVAMP. MathOctopusP 13B and ChatGPT are selected for comparison
6.2 LLM Gradually Learn to Align and Improve Multilingual Reasoning during Preference Optimization
To better analyze the state of the model during the optimization preference process, we tested the checkpoints of the DPO and PPO-LoRA training on MSVAMP and visualized the accuracy and ACR respectively in Figure 4.
The results show that the optimization of DPO is quite efficient. DPO enabled the MathOctopusP 7B model to achieve a higher accuracy than both ChatGPT and MathOctopusP 13B within just 100 steps. As the optimization continues, the accuracy remains relatively stable while the degree of alignment still shows an increasing trend.
On the other hand, although the optimization speed of PPO is slower, it took fewer than 1000 steps to surpass the larger-scale MathOctopusP 13B. As the model continued to explore and optimize preferences, it achieved superior multilingual reasoning performance compared to ChatGPT at 1600 steps.
The experimental results demonstrate that the model, through exploration and preference optimization, continuously aligns non-English reasoning with English reasoning, effectively enhancing its multilingual reasoning capabilities. For PPO-LoRA, although optimization was halted around 2700 steps due to computational limitations, it is observable that the growth in multilingual reasoning consistency and ability continues even beyond 2500 steps, showcasing the potential of our preference optimization framework.
6.3 Case Study
In the table example, we present an instance illustrating the impact of alignment on the model’s output. The baseline model, when presented with a question in English, correctly determined that an addition operation was required to solve for the suggested number of students to include bacon. However, when faced with the same question expressed in Chinese, it incorrectly assessed the variable computation relationship. This further demonstrates that although the original model was fine-tuned for multiple languages, there is an inconsistency in its internal reasoning capabilities and thought processes, leading to errors in non-English reasoning. After alignment, the model’s overall reasoning process became more similar to its English responses, and it correctly identified the numerical relationship. This example directly illustrates how our framework improves reasoning abilities by aligning the reasoning in other languages with English.
Figure 5: Accuracy across ten languages on MSVAMP after training MathOctopusP7B on preference datasets constructed using translation models of different scales.
6.4 Robustness over Different Scale of Translation Model
We also conducted ablation experiments comparing the translation models used during the preference estimation phase. In addition to NLLB-600M-distilled, we employed NLLB-1.3B-distilled  [^5] https://huggingface.co/facebook/nllb-200-distilled-1.3B and NLLB-3.3B  [^6] https://huggingface.co/facebook/nllb-200-3.3B, keeping the sampling dataset the same, and used them to estimate preferences, repeating the aforementioned experiments. The results are shown in Figure 5, where we found that all three models achieved impressive results, showing better reasoning capabilities over the Base Model and ChatGPT across ten languages. These results demonstrate that our method is robust for translation models of various sizes.
7 Conclustion
In this paper, we propose a novel multilingual alignment-as-preference optimization framework, significantly enhancing reasoning ability in non-pivot languages by aligning them with pivot languages. Specifically, we utilize an open-source translation model to estimate alignment preferences, which are then optimized using either DPO or PPO. Experimental results demonstrate that our framework achieved improvements in reasoning accuracy across three test sets, with a notable 13.7% increase in reasoning accuracy on out-of-domain datasets. The analysis also confirms that enhancing alignment through our method is the key to improvements in multilingual reasoning capabilities. Furthermore, we found that iterative DPO contributes to further alignment and enhancement of reasoning abilities, pushing the improvement to 16.7%.
Appendix A Appendix
A.1 Experiment Details
Our code is primarily based on the trl 777 https://github.com/huggingface/trl , with some minor modifications made. The modified code will also be made available at  [^8] https://github.com/NJUNLP/MAPO
Prompt: During the sampling, training, and testing phases, we consistently use the same prompt as MathOctopus Chen et al. (2023).
LoRA: For experiments using LoRA, such as DPO LoRA and PPO LoRA, we standardize the experimental settings, optimizing the [q_proj, v_proj , o_proj] modules. The values of r and alpha are set to 128 and 256, respectively.
DPO: We employ a learning rate of 1e-6, with $\beta$ set at 0.1, and a warmup step count of 100. The batch size is configured to 128. The optimization process is capped at a maximum of 1000 steps, and we save the checkpoint that corresponds to the lowest loss on validation set.
PPO: We have configured a learning rate of 2e-5, with a batch size of 64 and 2 PPO epochs. We utilize the AdamW optimizer to improve the stability of the optimization, setting the epsilon value to 1e-5. Additionally, we have implemented a linear learning rate warm-up technique for the first 150 steps. All other hyperparameters are set to the default values provided by the trl library. This is an appendix.